# Timer: Transformers for Time Series Analysis at Scale

### v1

@@ 4 Experiments
We demonstrate Timer as a large time series model in time series forecasting, imputation, and anomaly detection by tackling them in a unified generative scheme, which is detailed in the following sections. We also compare Timer with state-of-the-art approaches and present pre-training benefits on data-scarce scenarios, known as the few-shot ability of large models. Furthermore, we delve into the scalability, including the model size and data size. We also explore candidate backbones and architectures towards LTSMs, exhibiting the effectiveness of our architectural option. All downstream datasets evaluated in the experiment are not included in UTSD to prevent data leakage. We provide the detailed implementation and model configurations of pre-training and fine-tuning in Appendix B.1 and Appendix B.2.
@@ 4.1 Time Series Forecasting
Time series forecasting is essential and presents challenges in real-world applications. To thoroughly evaluate the performance, we elaborately establish the benchmark, including ETT, ECL, Traffic, Weather, and PEMS adopted in Liu et al. (2023). We adopt the unified lookback length as $672$ and the forecast length as $96$. We pre-training Timer on UTSD-12G with the segment length $S=96$ and the number of tokens $N=15$, such that Timer can deal with time series with the context length up to $1440$. And the downstream forecasting task is naturally converted into the next token prediction as detailed in Appendix B.2.
Figure 5: Downstream forecasting results of Timer obtained by training from scratch and fine-tuning from the pre-trained model. The detailed results are provided in Table 7.
@@ Results
As depicted in Figure 5, we present the results of the pre-trained Timer (solid line) and Timer trained from scratch (dashed line) under different data scarcities. Additionally, we also present previous state-of-the-art forecasters trained on full samples, that is, PatchTST (Nie et al., 2022) on ETTh1 and Weather; and iTransformer (Liu et al., 2023) on other datasets, as a competitive baseline. It is notable that the pre-trained Timer demonstrates superior performance to the baseline in four datasets with limited samples, specifically achieving better results with only $1\%$ available samples from ETTh1, $5\%$ from Traffic, $3\%$ from PEMS03, and $25\%$ from PEMS04 with remarkable few-shot ability.
Moreover, the pre-trained Timer exhibits impressive generalization capability through large-scale pre-training, which significantly outperforms the model trained from scratch. Concretely, the performance of training on the entire downstream dataset can be achieved by leveraging only $2\%$ of the training samples in ETTh1, $5\%$ in ECL, $1\%$ in Weather, and $4\%$ in PEMS03, which exemplifies the transferable knowledge acquired during pre-training on UTSD. Further promotions are achieved by training on all available samples in several undersaturated datasets: the forecasting error is reduced as $0.165\to 0.154$ on Weather, $0.126\to 0.118$ on PEMS03, and $0.125\to 0.107$ on PEMS04. Overall, the performance decline of pre-trained Timer is notably less pronounced with increased data scarcity, thereby alleviating the demand for downstream training samples.
@@ 4.2 Imputation
Imputation is widely required in real-world applications, aiming to fill the corrupted time series based on the observed data. However, while various machine learning algorithms and simple linear interpolation can effectively cope with the corruptions randomly happening at the point level, real-world corruptions typically result from prolonged monitor shutdowns and require a continuous period of recovery. Consequently, imputation can be ever challenging when attempting to recover a span of time points encompassing intricate series variations. In this task, we conduct the segment-level imputation. Each time series is divided into $8$ segments and each segment has the length of $24$ and the possibility of being completely masked. We obtain Timer on UTSD-4G by generative pre-training with the segment length $S=24$ and the token number $N=15$. For downstream adaptation, we conduct the denoising autoencoding in T5 (Raffel et al., 2020) as detailed in Appendix B.2 to recover the masked spans in an autoregressive way.
Figure 6: Pre-training benefit of Timer on $5\%$ available samples. We adopt mask ratios in $\{12.5\%,25\%,37.5\%,50\%\}$ and report the average results. Additional results are provided in Figure 14.
@@ Results
We conduct experiments on the same datasets of forecasting. As pre-training consistently yields positive effects on all datasets, we show the promotion as the reduction ratio of the imputation error in Figure 6. Additional results in Figure 14 also demonstrate remarkable error reduction with $20\%$ and $100\%$ available samples. We also evaluate previous state-of-the-art TimesNet (Wu et al., 2022) in Table 11 on the segment-wise imputation, where Timer outperforms TimesNet in respectively $100.0\%$, $86.4\%$, and $56.8\%$ of $44$ imputation settings under the same data-scarcities of $\{5\%,20\%,100\%\}$, validating the effectiveness of Timer for the challenging segment-level imputation.
@@ 4.3 Anomaly Detection
Anomaly detection is vital in industry and operations. Previous methods (Xu et al., 2021; Wu et al., 2022) generally adopt the reconstructive approach to tackle the unsupervised scenario, where a feature extractor is trained to reconstruct input series and the output is regarded as standard values. Based on our generative model, we cope with anomaly detection in a predictive approach. The core is to use the observed segments to predict the future segment, and the predicted segment will be established as the standard to be compared with the actual value received. Unlike previous methods requiring to collect time series of a period for reconstruction, our predictive approach allows for segment-level anomaly detection on the fly. Consequently, the task is converted into a forecasting task as detailed in Appendix B.2. The performance is evaluated on UCR Anomaly Archive (Wu & Keogh, 2021), where the task is to find the position of an anomaly in the test series based on a single normal series for training. We use the MSE between the predicted segment and ground truth as the confidence level.
@@
Results
We calculate the confidence of all segments on the test set and sort them. If the segment with the $\alpha$ quantile hits the anomaly position labeled in the dataset, it accomplishes the detection of the anomaly. Figure 7 compares the detection performance of pre-trained models and from scratch using two indicators. The left figure shows the number of datasets that the model has completed detection within the quantile of $3\%$ and $10\%$, while the right figure shows the quantile distribution and averaged quantile of all UCR datasets, where the pre-trained Timer with the smaller averaged quantile works as a more accurate detector.
@@ 4.4 Scalability
Scalability is the essential property that emerges from pre-trained models to large models. To investigate the scaling behavior of Timer, we conduct extensive pre-training with increased model size and data size as detailed in Table 4 and adapt for downstream forecasting on all subsets of PEMS.
@@
We keep UTSD-4G as the pre-training corpus. The results are depicted in the Figure 8. While maintaining a unified model dimension $D=256$, we increase the number of layers, yielding pre-trained models with parameters ranging from 1M to 4M. It leads to the reduction in forecasting errors of the two few-shot scenarios by an average of $14.7\%$ and $20.6\%$ respectively. Again, we continue to increase the model dimension under the layer number $L=6$ and enlarge the parameters from 3M to 50M, leading to further improved performance by $25.1\%$ and $18.2\%$, validating the effect of scaling up the model size.
@@
Data size
We pre-train Timer under the unified model size with different UTSD sizes, which exhibits steady improvement with the enlarged pre-training scale in Figure 9. The benefit is relatively small compared to expanding the model size from the beginning, which could be attributed to the performance saturation on this benchmark. Compared with large language models, the parameters of Timer can be still small, indicating the higher parameter efficiency in the time series modality, which is also supported by prior works (Das et al., 2023b). As the scaling law (Kaplan et al., 2020) of large models highlights the significance of synchronized scaling of data with the model parameters, there is still an urgent need to accelerate the data infrastructure in the time series field to promote the development of LTSMs.
Overall, by increasing the parameters and pre-training scale, Timer achieves $0.231\to 0.138\ (-40.3\%)$ and $0.194\to 0.123\ (-36.6\%)$ under few-shot scenarios, even surpassing the performance of the previous state-of-the-art (Liu et al., 2023) training on full samples of PEMS datasets $(0.139)$.
4.5 Analysis
Backbone for LTSM
We apply the unified model configuration and training strategy to pre-train the candidates on UTSD-4G. The loss curves of training and validation are provided in Figure 10, which confirms that Transformer exhibits excellent scalable ability as the backbone for large time series models, whereas MLP-based and CNN-based architectures may encounter the bottleneck in accommodating diverse time series data.
Decoder-only v.s. Encoder-only
While a smaller training loss is achieved by the encoder-only Transformer in Figure 10, the progress of large language models indicates that decoder-only models may possess stronger generalization capabilities in downstream adaptation (Wang et al., 2022a; Dai et al., 2022), which is the essential purpose of LTSMs. Therefore, we further compare the forecasting performance when adapted to different levels of data scarcity.
We elaborately evaluate the performance on six benchmarks in Table 1. By training from scratch, the encoder-only Transformer achieves better end-to-end performance in undersaturated scenarios (1% Target - None), while the decoder-only architecture requires more samples by training from scratch. However, after pre-training, Timer as the decoder-only Transformer exhibits significantly better generalization than the encoder-only pre-trained model, thereby enhancing the performance on most downstream scenarios. The observation partially elucidates why the encoder-only structure has become prevalent in the mainstream time series forecasting field; namely, the encoder-only model is more suitable for small benchmarks, while the decoder-only architecture, which is endowed notable generalization ability and model capacity, is a more suitable choice for developing LTSMs.
Table 1: Downstream forecasting results under different data scarcity of the encoder-only and decoder-only Transformer respectively pre-trained on UTST-12G. Datasets are ordered by the oversaturation in Figure 1. Full results of PEMS and ETT can be found in Table 14.

### v2

@@ 4 Experiments
We demonstrate Timer as a large time series model in time series forecasting, imputation, and anomaly detection by tackling them in a unified generative scheme, which is described in Figure 5. We compare Timer with state-of-the-art task-specific models and present the benefit of pre-training on data-scarce scenarios, known as the few-shot ability of large models. Furthermore, we delve into the model scalability, including model/data size, and try to build a comprehensive zero-shot evaluation across concurrent large time series models. All the downstream datasets are not included in the pre-training stage to prevent data leakage. We provide the detailed implementation and model configurations of pre-training and adaptation in Appendix B.1 and B.2.
Figure 5: Illustration of our generative task unification: (1) Generative pre-trained Timer can naturally predict the next series by the iterative autoregression; (2) By introducing masked tokens during adaptation, Timer generates imputations with the previous context and assemble them with the observed part; (3) We propose predictive anomaly detection by predicting normal series in advance.
@@ 4.1 Time Series Forecasting
Time series forecasting is essential and presents challenges in real-world applications. To thoroughly evaluate the performance, we elaborately establish the benchmark, including ETT, ECL, Traffic, Weather, and PEMS adopted in Liu et al. (2023b). We adopt the unified lookback length as $672$ and the forecast length as $96$. We pre-training Timer on UTSD-12G with the segment length $S=96$ and the number of tokens $N=15$, such that Timer can deal with time series with a context length up to $1440$. The downstream forecasting task can be naturally completed as the next token prediction, which is detailed in Appendix B.2.
Figure 6: Forecasting performance of Timer obtained by training from scratch and fine-tuning from the pre-trained model on different data scarcities. State-of-the-art small deep forecasters trained on full samples are provided the SOTA baseline. A smaller MSE indicates better results. Detailed results are provided in Table 10.
Figure 7: Performance comparison with state-of-the-art small deep models. For imputation, Time is compared with TimesNet (Wu et al., 2022) under different data scarcities, each of which contains $44$ imputation scenarios. In UCR Anomaly Detection Archive (Wu & Keogh, 2021), we compare the number of detected anomalies under given confidence quantiles. Detailed results are provided in Table 11-15.
@@ Results
As depicted in Figure 6, we present the results of the pre-trained Timer (solid line) and Timer trained from scratch (dashed line) under different data scarcities. We also evaluate state-of-the-art forecasters by training them on full samples as a competitive baseline. Concretely, we train PatchTST (Nie et al., 2022) and iTransformer (Liu et al., 2023b) on each dataset and report the better one as SOTA. Timer fine-tuned on few training samples demonstrates competitive results as advanced small deep forecasters, specifically achieving better results with only $1\%$ available samples from ETTh1, $5\%$ from Traffic, $3\%$ from PEMS03, and $25\%$ from PEMS04 and exhibiting remarkable few-shot ability.
To assess pre-training benefit, we compare solid and dashed lines, differing by whether to load the pre-trained checkpoint. Concretely, the performance of training a random-initialized Timer on full samples can be achieved by fine-tuning our pre-trained Timer with only $2\%$ of the training samples in ETTh1, $5\%$ in ECL, $1\%$ in Weather, and $4\%$ in PEMS03, which exemplifies the transferable knowledge acquired by pre-training on UTSD. When all samples are available, the performance of the pre-trained Timer can also outperform training it from scratch: the prediction error is reduced as $0.165\to 0.154$ on Weather, $0.126\to 0.118$ on PEMS03, and $0.125\to 0.107$ on PEMS04. Overall, in widespread data-scarce scenarios, the performance degradation can be alleviated by the few-shot generalization of LTSMs.
@@ 4.2 Imputation
Imputation is ubiquitous in real-world applications, aiming to fill corrupted time series based on partially observed data. However, while various machine learning algorithms and simple linear interpolation can effectively cope with the corruptions randomly happening at the point level, real-world corruptions typically result from prolonged monitor shutdowns and require a continuous period of recovery. Consequently, imputation can be ever challenging when attempting to recover a span of time points encompassing intricate series variations. In this task, we conduct the segment-level imputation. Each time series is divided into $8$ segments and each segment has the length of $24$ and the possibility of being completely masked. We obtain Timer on UTSD-4G by generative pre-training with the segment length $S=24$ and the token number $N=15$. For downstream adaptation, we conduct the denoising autoencoding in T5 (Raffel et al., 2020) as detailed in Appendix B.2 to recover the masked spans in a generative way.
Figure 8: Pre-training benefit of Timer on the downstream imputation task with $5\%$ available samples. Following TimesNet (Wu et al., 2022), each dataset is imputed with four mask ratios in $\{12.5\%,25\%,37.5\%,50\%\}$ and we calculate the average reduced imputation error in MSE relative to training from scratch. Additional results of other data scarcities are provided in Figure 18.
@@ Results
We establish a comprehensive segment-level imputation benchmark, which includes $11$ datasets with four mask ratios each. Timer is compared with the previous state-of-the-art imputation model (Wu et al., 2022). As shown in the left of Figure 7, Timer outperforms in respectively $100.0\%$, $86.4\%$, and $56.8\%$ of $44$ imputation scenarios under the data scarcities of $\{5\%,20\%,100\%\}$, validating the effectiveness of Timer on the challenging imputation task. Regarding the benefit of pre-training, we present the promotion as the reduction ratio of imputation errors in Figure 8, where pre-training consistently yields positive effects with $5\%$ downstream samples. Additional experiments on $20\%$ and $100\%$ available samples are provided in Figure 18, which still present notable performance improvement.
@@ 4.3 Anomaly Detection
Anomaly detection is vital in industry and operations. Previous methods (Xu et al., 2021; Wu et al., 2022) typically tackle the unsupervised scenario in a reconstructive approach, where a model is trained to reconstruct the input series, and the output is regarded as the normal series. Based on our generative model, we cope with anomaly detection in a predictive approach, which utilizes the observed segments to predict the future segment, and the predicted segment will be established as the standard to be compared with the actual value received. Unlike the previous method requiring to collect time series of a period for reconstruction, our predictive approach allows for segment-level anomaly detection on the fly. Thus, the task is converted into a next token prediction task as detailed in Appendix B.2.
We introduce UCR Anomaly Archive (Wu & Keogh, 2021) that contains $250$ tasks. In each task, a single normal time series is provided for training, and the model should locate the position of an anomaly in the test series. We first train a predictive model on the training set and calculate the MSE between the predicted series and ground truth on the test set. By regarding the MSE of all segments as the confidence level, the segments with higher than $\alpha$ quantile of confidence are labeled as potential positions of anomalies.
Results
We evaluate well-acknowledged anomaly detection models, including TimesNet (Wu et al., 2022) and Anomaly Transformer (Xu et al., 2021). As shown in the right of Figure 7, we present the number of detected anomalies with given quantiles, where Timer outperforms other advanced anomaly detection models, exhibiting the versatility of our generative time series model. Moreover, Figure 9 compares the detection performance of pre-trained models and from scratch using two indicators. The left figure shows the number of datasets that the model has completed detection within the quantile of $3\%$ and $10\%$, and the right figure shows the quantile distribution and the averaged quantile of all $250$ UCR datasets, where the pre-trained Timer with the smaller averaged quantile works as a more accurate detector.
@@ 4.4 Scalability
Scalability is the essential property that emerges from pre-trained models to large models. To investigate the scaling behavior of Timer, we pre-train Timer with increased model size and data size as detailed in Appendix B.1 and evaluate it in downstream forecasting on all subsets of PEMS.
@@
We keep UTSD-4G as the pre-training set. Results are presented in Figure 10. While keeping model dimension $D=256$, we increase the number of layers. The growth of parameters from 1M to 4M leads to the decrease in forecasting errors in two few-shot scenarios by an average of $14.7\%$ and $20.6\%$ respectively. Subsequently, we increase the model dimension under the fixed layer number $L=6$, enlarging parameters from 3M to 50M, resulting in further improved performance of $25.1\%$ and $18.2\%$, validating the efficacy of scaling up the model size.
@@
Data scale
We pre-train Timer under the same model size with different UTSD sizes, which exhibits steady improvement with the enlarged pre-training scale in Figure 11. The benefit is relatively small compared to expanding the model size previously, which can be attributed to the performance saturation on these datasets. Compared with large language models, the parameter scale of Timer can still be small, indicating the higher parameter efficiency in the time series modality, which is also supported by prior works (Das et al., 2023b). As the scaling law (Kaplan et al., 2020) of large models highlights the significance of synchronized scaling of data with the model parameters, there is still an urgent need to accelerate the data infrastructure in the time series field to promote the development of LTSMs.
Overall, by increasing the model size and data scale, Timer reduces the prediction error as $0.231\to 0.138\ (-40.3\%)$ and $0.194\to 0.123\ (-36.6\%)$ under few-shot scenarios, surpassing state-of-the-art multivariate forecaster (Liu et al., 2023b) training on full samples of PEMS datasets $(0.139)$.
Table 1: Downstream forecasting results under different data scarcity of the encoder-only and decoder-only Transformer respectively pre-trained on UTST-12G. Datasets are ordered by the degradation in Figure 1. Full results of PEMS and ETT can be found in Table 17.